a motorboat covers a certain downstream in a river in 5 hours it covers the same distance upstream in five hour and a half the speed of the water is 1.5 km. find the speed of the boat in still water.
Answer by edjones(8007)

d=ts d=distance t=time, s=speed
ts=ts
5(s+1.5)=5.5(s-1.5)
5s+7.5=5.5s-8.25
15.75=.5s add -5s+8.25 to each side.
divide each side by .5
s=31.5 km/h
Always check:
5*33=165 km
5.5*30=165 km
Ed